OpenStudy (anonymous):

Rationalize the denominator of sqrt(5/13) and sqrt(x/6) and sqrt(y/2z)

OpenStudy (anonymous):

these are the problems I need to rationalize the denominator for \[\sqrt(5/13)\] also \[\sqrt(x/6)\] also \[\sqrt(y/2z)\]

OpenStudy (anonymous):

Just Multiply each one of them by the \[\sqrt{Denomin}\] both the Top and the bottom should be multiplied

OpenStudy (anonymous):

Example the first shld be multiplied by \[\sqrt{13}\] BOTH ABOVE AND BELOW THE FRACTION LINE

OpenStudy (anonymous):

then ur denom becomes rational = 13

OpenStudy (anonymous):

Use of course this \[\sqrt{A}*\sqrt{B} = \sqrt{AB}\]

OpenStudy (unklerhaukus):

\[\sqrt\frac 5{13}=\sqrt{\frac 5{13}}\times 1\]\[\qquad\qquad=\sqrt\frac 5{13}\times\sqrt\frac{13}{13}\]\[\qquad\qquad=\sqrt\frac 5{13}\times\sqrt\frac{13}{13}\]\[\qquad\qquad=\frac{\sqrt{5}\sqrt{13}}{13}\]

OpenStudy (anonymous):

This is NOT COMPLETE to the end - it should be written

OpenStudy (anonymous):

\[\frac{ \sqrt{65} }{ 13 }\]

OpenStudy (anonymous):

SPOTLESS HAVE U GOT MY IDEA ?

OpenStudy (anonymous):

O.K. good and farewell :)

OpenStudy (anonymous):

Yes Mikael, once again it's very much appreciated. I found a youtube tutorial to better explain the concept you've just showed me. Appreciate the help!

OpenStudy (anonymous):

Glad to make ur life easier a bit :)